OpenStudy (anonymous):

Solve dy/dx=(2cos^2(x)-sin^2(x)+y^2)/2cos(x) given that y=sin(x)+1/v(x)

OpenStudy (anonymous):

\[dy/dx = \left( 2\cos ^{2}x -\sin ^{2}x+y ^{2} \right) / 2cosx\]

OpenStudy (anonymous):

Need to solve this equation via general substitution given that y = sinx + 1/v(x)